# High-dimensional sliced inverse regression with endogeneity

### v1

@@ 5 Simulation studies
We conducted a numerical study to evaluate the performance of the proposed 2SLSIR estimator for recovering the (sparsity patterns of the) central subspace. We compared this against two alternatives: the sparse Lasso SIR (LSIR) estimator of Lin et al., (2018, 2019) which ignores endogeneity and does not use the instruments, and, in the case of single index model $(d=1)$, the two-stage Lasso (2SLasso) estimator of Lin et al., (2015) which assumes linearity and additive errors in the response model i.e., $y_{i}=\mathbf{X}_{i}^{\top}\mathbf{B}+\varepsilon_{i}$.
@@
We simulated data $\{(y_{i},\mathbf{X}_{i},\mathbf{Z}_{i});i=1,\ldots,n\}$, from a linear instrumental variable model $\mathbf{X}_{i}=\bm{\Gamma}^{\top}\mathbf{Z}_{i}+\mathbf{U}_{i}$ coupled with one of the following four models for the response:
$(i):\;y_{i}$ $=\exp(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i})\quad(ii):\;y_{i}=(1/2)(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i})^{3}$
$(iii):\;y_{i}$ $=(\mathbf{X}_{i}^{\top}\bm{\beta}_{2})\exp(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}/4+\varepsilon_{i})\quad(iv):\;y_{i}=\frac{\exp(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i})}{3/2+\mathbf{X}_{i}^{\top}\bm{\beta}_{2}+\varepsilon_{i}}.$
These single and double index models are similar to those previously used in the literature to assess SDR methods (e.g., Lin et al.,, 2019). The true sparse matrix $\mathbf{B}=\bm{\beta}_{1}$ in the single index models (i) and (ii), and $\mathbf{B}=[\bm{\beta}_{1},\bm{\beta}_{2}]$ in the multiple index models (iii) and (iv), was constructed by randomly selecting $s=5$ nonzero rows of $\mathbf{B}$ and generating each entry in these rows from a uniform distribution on the two disjoint intervals $[-1,-0.5]$ and $[0.5,1]$, which we denote as $U([-1,-0.5]\cup[0.5,1])$. All remaining entries are set to zero. We assumed the number of dimensions $d$ was known throughout the simulation study.
For the instrumental variable model, we followed the same generation mechanism as in Lin et al., (2015). Specifically, to represent different levels of instrument strength, we considered three choices for $\bm{\Gamma}$: i) all instruments are strong, where we sampled $r=5$ nonzero entries for each column from the uniform distribution $U([-b,-a]\cup[a,b])$ with $(a,b)=(0.75,1)$; ii) all instruments are weak, where we sampled the same number of nonzero entries but with $(a,b)=(0.5,0.75)$; iii) instruments are of mixed strength, such that for each column of $\bm{\Gamma}$ we sampled $r=50$ nonzero entries comprising five strong/weak instruments with $(a,b)=(0.5,1)$ and 45 very weak instruments with $(a,b)=(0.05,0.1)$. Next we generated the random noise vector $(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}$ from a $(p+1)$-variate Gaussian distribution with zero mean vector and covariance matrix $\bm{\Sigma}=(\sigma_{ij})$ specified as follows. We set $\sigma_{ij}=(0.2)^{|i-j|}$ for $i,j=$ $1,\ldots,p$, and $\sigma_{p+1,p+1}=1$. For the last column/row of $\bm{\Sigma}$ characterizing endogeneity, the nonzero entries were then given by the nonzero row of $\mathbf{B}$, with five of these entries randomly set to $0.3$. Finally, we generated the instruments $\mathbf{Z}_{i}$ from a $q$-variate Gaussian distribution with zero mean vector and two choices for the covariance matrix $\bm{\Sigma}_{\mathbf{Z}}$, namely the $q\times q$ identity matrix and the AR(1) covariance matrix with autocorrelation $0.5$. In Supplementary Material S4, we conducted additional simulation results where the instruments $\mathbf{Z}$ were generated from a Bernoulli distribution, with largely similar trends to those seen below.
We set $p=q=100$ and varied the sample size $n\in\{500,1000\}$, simulating $100$ datasets for each configuration. For each simulated dataset, all tuning parameters across the three methods (2SLSIR, LSIR, 2SLasso) were selected via ten-fold cross-validation. We assessed performance using the estimation error on the projection matrix $\|\mathcal{P}(\mathbf{B})-\mathcal{P}(\hat{\mathbf{B}})\|_{F}$, along with the F1 score for variable selection.
Across both single and double index models (Tables 1 and 2, respectively), the proposed 2SLSIR estimator performed best overall in terms of both estimation error and variable selection. Indeed, it is the only estimator among the three whose estimation error consistently decreased when sample size increased from $n=500$ to $n=1000$; this lends empirical support to estimation consistency results established in Section 4. For the single index models (i) and (ii), the two-stage Lasso (2SLasso) estimator performed even worse than the single stage Lasso SIR (LSIR) estimator in terms of estimation error, particularly when the degree of non-linearity in the response model was relatively large.
@@
Table 1: Simulation results for the Lasso SIR (LSIR) estimator, the two-stage Lasso (2SLasso) estimator, and the proposed two-stage Lasso SIR (2SLSIR) estimator, in the single index models (i) and (ii). Standard errors are included in parentheses. The lowest estimation error and the highest F1 score are highlighted in each row.
$\bm{\Sigma}_{\mathbf{Z}}$ $n$ Model $\bm{\Gamma}$ Error F1
LSIR 2SLasso 2SLSIR LSIR 2SLasso 2SLSIR
Diag 500 (i) Strong 0.16 (0.05) 1.06 (0.22) 0.12 (0.05) 1.00 (0.01) 0.51 (0.18) 1.00 (0.00)
Weak 0.26 (0.07) 1.05 (0.21) 0.17 (0.08) 0.99 (0.03) 0.51 (0.19) 1.00 (0.01)
Mixed 0.16 (0.05) 1.06 (0.22) 0.13 (0.06) 1.00 (0.00) 0.54 (0.19) 1.00 (0.00)
(ii) Strong 0.16 (0.05) 0.28 (0.09) 0.12 (0.05) 1.00 (0.01) 0.96 (0.07) 1.00 (0.00)
Weak 0.26 (0.06) 0.33 (0.10) 0.19 (0.07) 0.99 (0.03) 0.95 (0.08) 1.00 (0.01)
Mixed 0.16 (0.06) 0.25 (0.08) 0.12 (0.05) 1.00 (0.00) 0.98 (0.04) 1.00 (0.00)
1000 (i) Strong 0.17 (0.05) 0.97 (0.22) 0.09 (0.04) 1.00 (0.00) 0.57 (0.18) 1.00 (0.00)
Weak 0.28 (0.06) 0.93 (0.23) 0.13 (0.05) 0.99 (0.03) 0.58 (0.20) 1.00 (0.00)
Mixed 0.20 (0.05) 0.96 (0.22) 0.09 (0.03) 1.00 (0.01) 0.58 (0.15) 1.00 (0.00)
(ii) Strong 0.18 (0.06) 0.18 (0.07) 0.09 (0.03) 1.00 (0.00) 1.00 (0.02) 1.00 (0.00)
Weak 0.29 (0.07) 0.22 (0.08) 0.12 (0.05) 0.99 (0.04) 0.99 (0.04) 1.00 (0.00)
Mixed 0.20 (0.06) 0.17 (0.05) 0.08 (0.03) 0.99 (0.04) 1.00 (0.01) 1.00 (0.00)
AR1 500 (i) Strong 0.18 (0.06) 1.09 (0.19) 0.13 (0.06) 1.00 (0.00) 0.53 (0.18) 1.00 (0.00)
Weak 0.29 (0.08) 1.07 (0.23) 0.19 (0.09) 0.99 (0.04) 0.56 (0.21) 1.00 (0.01)
Mixed 0.19 (0.06) 1.07 (0.18) 0.14 (0.06) 1.00 (0.01) 0.52 (0.14) 1.00 (0.00)
(ii) Strong 0.18 (0.06) 0.28 (0.09) 0.14 (0.05) 1.00 (0.01) 0.97 (0.06) 1.00 (0.00)
Weak 0.29 (0.06) 0.34 (0.11) 0.20 (0.08) 0.99 (0.04) 0.94 (0.08) 1.00 (0.02)
Mixed 0.18 (0.07) 0.27 (0.09) 0.13 (0.06) 1.00 (0.03) 0.97 (0.06) 1.00 (0.00)
1000 (i) Strong 0.21 (0.07) 1.03 (0.21) 0.10 (0.04) 1.00 (0.02) 0.55 (0.17) 1.00 (0.00)
Weak 0.33 (0.07) 0.99 (0.23) 0.14 (0.06) 0.97 (0.05) 0.57 (0.19) 1.00 (0.00)
Mixed 0.25 (0.06) 1.04 (0.21) 0.10 (0.04) 0.99 (0.02) 0.54 (0.17) 1.00 (0.00)
(ii) Strong 0.22 (0.06) 0.19 (0.06) 0.09 (0.04) 1.00 (0.03) 0.99 (0.03) 1.00 (0.00)
Weak 0.33 (0.07) 0.23 (0.07) 0.13 (0.06) 0.98 (0.06) 0.98 (0.04) 1.00 (0.00)
Mixed 0.25 (0.06) 0.18 (0.05) 0.09 (0.04) 0.99 (0.03) 1.00 (0.02) 1.00 (0.00)
@@
Table 2: Simulation results for the Lasso SIR (LSIR) estimator and the proposed two-stage Lasso SIR (2SLSIR) estimator, in the double index models (iii) and (iv). Standard errors are included in parentheses. The lowest estimation error and the highest F1 score are highlighted in each row.
$\bm{\Sigma}_{\mathbf{Z}}$ $n$ Model $\bm{\Gamma}$ Error F1
LSIR 2SLSIR LSIR 2SLSIR
diag 500 (iii) Strong 0.81 (0.22) 0.67 (0.24) 0.64 (0.14) 0.76 (0.13)
Weak 1.00 (0.21) 0.98 (0.24) 0.58 (0.13) 0.67 (0.16)
Mixed 0.92 (0.22) 0.71 (0.24) 0.52 (0.14) 0.77 (0.15)
(iv) Strong 0.42 (0.16) 0.41 (0.17) 0.92 (0.11) 0.91 (0.11)
Weak 0.51 (0.19) 0.49 (0.21) 0.86 (0.15) 0.86 (0.13)
Mixed 0.44 (0.20) 0.43 (0.21) 0.89 (0.12) 0.90 (0.12)
1000 (iii) Strong 0.91 (0.21) 0.47 (0.21) 0.58 (0.14) 0.87 (0.14)
Weak 1.04 (0.20) 0.68 (0.27) 0.55 (0.11) 0.78 (0.14)
Mixed 0.99 (0.16) 0.50 (0.20) 0.47 (0.07) 0.86 (0.12)
(iv) Strong 0.36 (0.17) 0.29 (0.17) 0.95 (0.12) 0.97 (0.08)
Weak 0.46 (0.17) 0.34 (0.19) 0.88 (0.13) 0.94 (0.11)
Mixed 0.41 (0.20) 0.27 (0.15) 0.89 (0.15) 0.97 (0.08)
AR1 500 (iii) Strong 0.88 (0.25) 0.73 (0.28) 0.63 (0.15) 0.76 (0.15)
Weak 1.03 (0.22) 1.02 (0.30) 0.57 (0.13) 0.70 (0.17)
Mixed 0.99 (0.21) 0.76 (0.27) 0.51 (0.15) 0.75 (0.15)
(iv) Strong 0.46 (0.18) 0.44 (0.20) 0.89 (0.13) 0.89 (0.12)
Weak 0.55 (0.19) 0.53 (0.22) 0.84 (0.15) 0.83 (0.14)
Mixed 0.47 (0.22) 0.45 (0.24) 0.86 (0.14) 0.88 (0.13)
1000 (iii) Strong 0.99 (0.20) 0.52 (0.20) 0.54 (0.13) 0.84 (0.12)
Weak 1.09 (0.20) 0.70 (0.26) 0.52 (0.10) 0.78 (0.16)
Mixed 1.06 (0.16) 0.52 (0.24) 0.45 (0.07) 0.85 (0.14)
(iv) Strong 0.39 (0.16) 0.30 (0.18) 0.93 (0.10) 0.97 (0.08)
Weak 0.52 (0.17) 0.36 (0.18) 0.85 (0.12) 0.93 (0.11)
Mixed 0.46 (0.21) 0.31 (0.20) 0.86 (0.14) 0.96 (0.08)

### v2

@@ 5 Simulation studies
We conducted a numerical study to evaluate the performance of the proposed two-stage lasso SIR estimator for recovering the (sparsity patterns of the) dimension reduction subspace.We simulated data $\{(y_{i},\mathbf{X}_{i},\mathbf{Z}_{i});i=1,\ldots,n\}$, from a linear instrumental variable model $\mathbf{X}_{i}=\bm{\Gamma}^{\top}\mathbf{Z}_{i}+\mathbf{U}_{i}$ coupled with one of the following five models for the response:
$\text{(i)}:\;y_{i}$ $=\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i},~~\text{(ii)}:\;y_{i}=\exp(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i}),\quad\text{(iii)}:\;y_{i}=\sinh(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i})$
$\text{(iv)}:\;y_{i}$ $=(\mathbf{X}_{i}^{\top}\bm{\beta}_{2})\exp(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i}),\quad\text{(v)}:\;y_{i}=\frac{\exp(\mathbf{X}_{i}^{\top}\bm{\beta}_{1}+\varepsilon_{i})}{3/2+\mathbf{X}_{i}^{\top}\bm{\beta}_{2}+\varepsilon_{i}}.$
Model (i) is the linear model, while the single and double index models (ii) - (v) are similar to those previously used in the literature to assess SDR methods (e.g., Lin et al.,, 2019). The true sparse matrix $\mathbf{B}=\bm{\beta}_{1}$ in the single index models (i) and (ii), and $\mathbf{B}=[\bm{\beta}_{1},\bm{\beta}_{2}]$ in the multiple index models (iii) and (iv) are constructed as follows: We randomly selected $s=5$ rows of $\mathbf{B}$ to be nonzero and generated each non-zero entry in these rows from a uniform distribution on the two disjoint intervals [1,0.5][-1,-0.5] and $[0.5,1]$, which we denote as $U([-1,-0.5]\cup[0.5,1])$. The elements in all other rows were set to zero. We then used a Gram-Schmidt process to ensure $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ were orthogonal to each other in the case of the double index model. The covariate $\mathbf{X}_{i}$ vector was generated from the instrumental variable model $\mathbf{X}_{i}=\bm{\Gamma}^{\top}\mathbf{Z}_{i}+\mathbf{U}_{i}$ for $i=1,\ldots,n$, where we followed a similar generation mechanism to that in Lin et al., (2015). That is, we simulated each of the $q$ elements of $\mathbf{Z}_{i}$ from either: (1) a standard normal distribution, or (2) a Bernoulli distribution with success probability $0.5$. To generate $\bm{\Gamma}$, we randomly set $r=5$ nonzero entries in each column from the uniform distribution $U([-b,-a]\cup[a,b])$ with $(a,b)=(0.75,1)$. Finally, we generated the random noise vector $(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}$ from a $(p+1)$-variate Gaussian distribution with zero mean vector and covariance matrix $\bm{\Sigma}=(\sigma_{ij})$ as follows: We set $\sigma_{ij}=(0.2)^{|i-j|}$ for $i,j=$ $1,\ldots,p$, so that $\bm{\Sigma}_{\mathbf{U}}$ has an AR(1) structure. For the last column/row of $\bm{\Sigma}$ characterizing endogeneity, we set $\bm{\Sigma}_{S,p+1}=-\bm{\Sigma}_{S,S}\bm{\beta}_{1S}$, where $S$ denotes the index set of non-zero elements in $\bm{\beta}_{1}$, and randomly selected another five elements and set them equal to $0.3$, while the remaining elements were set to zero. The last element $\sigma_{p+1,p+1}=\bm{\Sigma}_{\mathbf{U},\varepsilon}^{\top}\bm{\Sigma}_{\mathbf{U},\mathbf{U}}\bm{\Sigma}_{\mathbf{U},\varepsilon}+U(0,0.2)$. This construction of $\bm{\Sigma}$ ensures it is positive definite, and that $\bm{\Sigma}_{\mathbf{X}}^{-1}\bm{\Sigma}_{\mathbf{X},\epsilon}$ does not belong to the column space of $\mathbf{B}$.
We considered both low-dimensional and high-dimensional settings, where for the former we set $p=q=40$ and varied $n\in\{200,500\}$, while for high-dimensional settings we fixed $n=200$ and varied $p=q\in\{500,1000\}$. For each configuration, we simulated $100$ datasets, where in each dataset we applied the one-stage lasso, lasso SIR, two-stage lasso, and the proposed two-stage lasso SIR estimators. Note both the one-stage lasso and two-stage lasso estimates assume $d=1$, while the dimension $d$ is assumed to be known for both the lasso SIR and two-stage lasso SIR. For the latter two SIR-based estimators, we also set the number of slices $H=10$. To reduce the computational cost, for the two-stage lasso and two-stage lasso SIR estimators, we selected the tuning parameter for the penalized regression of each covariate on instruments to minimize the Bayesian information criterion (BIC). We also found that using other information criterion, such as extended BIC of chen2008extended, led to similar results in the second stage in both low and high-dimensional settings. All other tuning parameters were selected via an appropriate ten-fold cross-validation approach. We assessed performance of all methods based on the error on the estimated projection matrix $\|\mathcal{P}(\hat{\mathbf{B}})-\mathcal{P}(\mathbf{B})\|_{F}$, along with the area under the receiver operating curve (AUC) score for variable selection performance.
Table 1 demonstrates that in low-dimensional settings, the proposed two-stage lasso SIR estimator performed best in terms of both estimation error and variable selection. Indeed, it is the only method among the four considered whose estimation error consistently decreased when sample size increased from $n=200$ to $n=500$. The one-stage lasso estimator, which ignores both endogeneity and non-linearity in the outcome model, performed the worst as expected. Comparing the lasso SIR and two-stage lasso, the former tended to perform better except in model (i), suggesting the consequence of ignoring non-linearity and/or underestimating the dimensions were more serious than ignoring endogeneity in this simulation design. Turning to the variable selection results, the lasso SIR estimator has overall good performance, while the two-stage lasso estimator only achieves an AUC close to 11 in the linear model (i). The proposed two-stage lasso SIR estimator consistently had the (equal) best selection performance.
For high-dimensional settings, Table 2 shows that the proposed two-stage lasso SIR still had the best performance overall. For the linear model (i), the two-stage lasso performed well as expected, but its performance was worse than the two-stage lasso SIR estimator in single index models (ii)-(iii) where the link function is non-linear. The proposed two-stage lasso SIR also had a much smaller estimation error compared with the one-stage lasso SIR in all single index models (i)–(iii), although the two methods had quite similar estimation error in the double index models (iv)-(v). Finally, the performance of all methods in both low- and high-dimensional settings was similar for both continuous and binary cases of $\mathbf{Z}$. In particular, this suggested the proposed two-stage lasso SIR estimator was not overly sensitive to violation of Condition 2 in this simulation design.
$\mathbf{Z}$ Model $n$ Error AUC
lasso LSIR 2Slasso 2SLSIR lasso LSIR 2Slasso 2SLSIR
Cont. (i) 200 0.62 (0.16) 0.47 (0.13) 0.25 (0.15) 0.18 (0.07) 0.99 (0.02) 0.99 (0.03) 1.00 (0.02) 1.00 (0.00)
500 0.65 (0.13) 0.44 (0.13) 0.14 (0.07) 0.09 (0.03) 0.99 (0.02) 1.00 (0.01) 1.00 (0.00) 1.00 (0.00)
(ii) 200 1.13 (0.17) 0.47 (0.13) 1.03 (0.20) 0.18 (0.07) 0.73 (0.12) 0.99 (0.03) 0.77 (0.13) 1.00 (0.00)
500 1.13 (0.14) 0.44 (0.13) 1.02 (0.17) 0.09 (0.03) 0.72 (0.12) 1.00 (0.01) 0.77 (0.14) 1.00 (0.00)
(iii) 200 1.10 (0.19) 0.47 (0.13) 0.95 (0.25) 0.18 (0.07) 0.73 (0.14) 0.99 (0.03) 0.81 (0.15) 1.00 (0.00)
500 1.05 (0.18) 0.44 (0.13) 0.90 (0.23) 0.09 (0.03) 0.78 (0.13) 1.00 (0.01) 0.83 (0.14) 1.00 (0.00)
(iv) 200 1.48 (0.24) 0.87 (0.19) 1.45 (0.24) 0.67 (0.22) 0.60 (0.10) 1.00 (0.01) 0.65 (0.09) 1.00 (0.01)
500 1.44 (0.22) 0.81 (0.21) 1.36 (0.22) 0.39 (0.17) 0.62 (0.09) 1.00 (0.01) 0.65 (0.13) 1.00 (0.00)
(v) 200 1.63 (0.12) 0.85 (0.21) 1.57 (0.15) 0.70 (0.21) 0.56 (0.07) 0.99 (0.02) 0.60 (0.11) 1.00 (0.01)
500 1.60 (0.19) 0.75 (0.22) 1.51 (0.23) 0.41 (0.19) 0.53 (0.06) 1.00 (0.01) 0.60 (0.07) 1.00 (0.02)
Bin (i) 200 0.61 (0.16) 0.46 (0.15) 0.25 (0.18) 0.17 (0.06) 0.98 (0.03) 0.99 (0.03) 1.00 (0.02) 1.00 (0.00)
500 0.68 (0.14) 0.48 (0.15) 0.14 (0.08) 0.09 (0.03) 0.99 (0.03) 0.99 (0.02) 1.00 (0.00) 1.00 (0.00)
(ii) 200 1.09 (0.18) 0.46 (0.15) 1.04 (0.21) 0.17 (0.06) 0.73 (0.11) 0.99 (0.03) 0.77 (0.12) 1.00 (0.00)
500 1.03 (0.21) 0.48 (0.15) 0.94 (0.22) 0.09 (0.03) 0.79 (0.13) 0.99 (0.02) 0.81 (0.13) 1.00 (0.00)
(iii) 200 1.07 (0.17) 0.61 (0.17) 0.94 (0.23) 0.22 (0.12) 0.76 (0.12) 0.98 (0.03) 0.81 (0.13) 1.00 (0.01)
500 0.99 (0.19) 0.67 (0.14) 0.82 (0.24) 0.13 (0.05) 0.82 (0.12) 0.98 (0.03) 0.88 (0.13) 1.00 (0.00)
(iv) 200 1.47 (0.22) 0.80 (0.22) 1.41 (0.19) 0.61 (0.22) 0.62 (0.09) 1.00 (0.01) 0.62 (0.07) 1.00 (0.01)
500 1.35 (0.21) 0.80 (0.22) 1.35 (0.20) 0.38 (0.17) 0.64 (0.07) 1.00 (0.01) 0.66 (0.09) 1.00 (0.00)
(v) 200 1.53 (0.18) 0.82 (0.20) 1.55 (0.15) 0.64 (0.20) 0.61 (0.10) 1.00 (0.01) 0.63 (0.12) 1.00 (0.01)
500 1.46 (0.23) 0.81 (0.22) 1.52 (0.19) 0.39 (0.16) 0.66 (0.14) 1.00 (0.02) 0.68 (0.16) 1.00 (0.00)
@@
Table 1: Simulation results for the lasso, one-stage lasso SIR (LSIR), two-stage lasso (2Slasso), and the two-stage lasso SIR (2SLSIR) estimators, in low-dimensional settings with $p=q=40$. Standard errors are included in parentheses. The lowest estimation error and the highest AUC score are highlighted in each row.
$\mathbf{Z}$ Model $p=q$ Error AUC
lasso LSIR 2Slasso 2SLSIR lasso LSIR 2Slasso 2SLSIR
Cont. (i) 500 0.31 (0.09) 0.41 (0.14) 0.13 (0.03) 0.25 (0.12) 1.00 (0.00) 0.97 (0.05) 1.00 (0.00) 0.98 (0.04)
1000 0.31 (0.10) 0.61 (0.16) 0.13 (0.04) 0.35 (0.13) 1.00 (0.00) 0.97 (0.05) 1.00 (0.00) 0.98 (0.04)
(ii) 500 1.27 (0.13) 0.39 (0.13) 1.22 (0.14) 0.21 (0.10) 0.66 (0.12) 0.99 (0.03) 0.72 (0.13) 0.99 (0.03)
1000 1.30 (0.14) 0.63 (0.16) 1.25 (0.14) 0.38 (0.14) 0.57 (0.06) 0.97 (0.05) 0.60 (0.07) 0.97 (0.05)
(iii) 500 1.16 (0.21) 0.39 (0.13) 1.11 (0.18) 0.21 (0.10) 0.70 (0.13) 0.98 (0.04) 0.73 (0.14) 0.99 (0.03)
1000 1.20 (0.15) 0.50 (0.18) 1.18 (0.17) 0.30 (0.15) 0.68 (0.12) 0.97 (0.06) 0.68 (0.13) 0.98 (0.04)
(iv) 500 1.65 (0.12) 1.05 (0.24) 1.64 (0.18) 1.00 (0.28) 0.55 (0.05) 0.99 (0.02) 0.53 (0.05) 1.00 (0.02)
1000 1.61 (0.22) 0.95 (0.26) 1.64 (0.17) 0.95 (0.33) 0.53 (0.05) 1.00 (0.01) 0.53 (0.05) 1.00 (0.02)
(v) 500 1.68 (0.08) 1.08 (0.25) 1.66 (0.11) 1.07 (0.28) 0.54 (0.07) 1.00 (0.02) 0.57 (0.10) 1.00 (0.02)
1000 1.69 (0.09) 1.02 (0.28) 1.73 (0.00) 1.02 (0.30) 0.52 (0.05) 0.99 (0.03) 0.50 (0.00) 0.99 (0.04)
Bin (i) 500 0.32 (0.09) 0.39 (0.15) 0.13 (0.04) 0.23 (0.11) 1.00 (0.00) 0.98 (0.05) 1.00 (0.00) 0.99 (0.04)
1000 0.31 (0.10) 0.47 (0.19) 0.13 (0.04) 0.28 (0.13) 1.00 (0.00) 0.97 (0.07) 1.00 (0.00) 0.98 (0.04)
(ii) 500 1.22 (0.19) 0.42 (0.13) 1.17 (0.16) 0.21 (0.09) 0.67 (0.13) 0.98 (0.04) 0.70 (0.11) 0.99 (0.03)
1000 1.26 (0.14) 0.45 (0.14) 1.23 (0.12) 0.29 (0.17) 0.62 (0.10) 0.97 (0.05) 0.68 (0.12) 0.97 (0.05)
(iii) 500 1.18 (0.16) 0.41 (0.14) 1.07 (0.18) 0.23 (0.11) 0.70 (0.12) 0.98 (0.04) 0.77 (0.13) 0.99 (0.03)
1000 1.19 (0.17) 0.48 (0.16) 1.13 (0.17) 0.30 (0.14) 0.70 (0.12) 0.97 (0.05) 0.74 (0.14) 0.97 (0.05)
(iv) 500 1.63 (0.19) 1.03 (0.24) 1.59 (0.16) 0.98 (0.24) 0.55 (0.08) 1.00 (0.02) 0.57 (0.08) 1.00 (0.01)
1000 1.61 (0.17) 0.94 (0.29) 1.67 (0.13) 0.94 (0.31) 0.54 (0.05) 0.99 (0.04) 0.53 (0.05) 0.99 (0.03)
(v) 500 1.63 (0.21) 1.08 (0.23) 1.65 (0.15) 1.05 (0.26) 0.56 (0.12) 0.99 (0.02) 0.62 (0.14) 1.00 (0.01)
1000 1.71 (0.04) 1.06 (0.28) 1.71 (0.06) 1.05 (0.28) 0.52 (0.04) 0.99 (0.04) 0.54 (0.10) 0.99 (0.03)
@@
Table 2: Simulation results for the lasso, lasso SIR (LSIR), the two-stage lasso (2Slasso), and the two-stage lasso SIR (2SLSIR) estimators, in high-dimensional settings with $n=200$. Standard errors are included in parentheses. The lowest estimation error and the highest AUC score are highlighted in each row.
6 Selection of dimensions
An important problem in applying any sufficient dimension reduction method in practice is to estimate the dimension $d$ for the dimension reduction subspace. When there is no endogeneity, this can be estimated by examining the eigenvalues of $\widehat{\bm{\Lambda}}^{*}$, which is an estimate of $\bm{\Lambda}^{*}=\operatorname{\text{cov}}\{E(\mathbf{X}\mid y)\}$. In classical settings with $p$ fixed and $n\to\infty$, it can be shown that the $d$ largest estimated eigenvalues of $\widehat{\bm{\Lambda}}^{*}$ are separated from the last HdH-d eigenvalues, and hence we can estimate $d$ via an asymptotic test (shao2007marginal) or a permutation/bootstrap test (e.g., ye2003using). The problem is more challenging in high-dimensional settings where $p/n\rightarrow 0$, since all non-zero eigenvalues of $\widehat{\bm{\Lambda}}^{*}$ are of the same order $p/n$ and no longer separated. To overcome this issue, Lin et al., (2019) proposed selecting the structural dimension for the lasso SIR estimator by clustering the adjusted eigenvalues $\tilde{\lambda}^{*}_{k}=\hat{\lambda}_{k}^{*}\|\hat{\bm{\beta}}_{k}^{*}\|_{2}$, where $\hat{\bm{\beta}}_{k}^{*}$ is a solution of (3) for $k=1,\ldots,H$. Other sparse SIR methods, such as those proposed by Tan et al., (2018); zeng2024subspace, select the structural dimension $d$ by augmenting an objective function with a nuclear norm constraint.
In our setting where endogeneity is present, the selection of dimensions in high dimensions is made more challenging by the complex eigenstructure of $\bm{\Lambda}=\operatorname{\text{cov}}\{E(\widehat{\mathbf{X}}\mid y)\}=\bm{\Gamma}^{\top}\operatorname{\text{cov}}\{E(\mathbf{Z}\mid y)\}\bm{\Gamma}$. A first natural solution to this problem is to apply the clustering method in Lin et al., (2019) in Stage 2 of the proposed two-stage estimator, when we compute the lasso SIR estimator of the outcome data $\mathbf{y}$ on the first stage fitted value matrix $\widehat{\mathcal{X}}=\mathcal{Z}\widehat{\bm{\Gamma}}$. On the other hand, we can also write (2) as $y_{i}=f(\mathbf{Z}_{i}^{\top}\mathbf{C},\tilde{\varepsilon}_{i})$, where $\mathbf{C}=\bm{\Gamma}\mathbf{B}$ is a $q\times d$ matrix. It follows that the central subspace of $y$ on $\mathbf{Z}$ also has dimension $d$, assuming $\mathbf{C}$ is full rank. Therefore, we also consider selecting $d$ simply by computing a lasso SIR estimator of the outcome $y$ on the instrumental variables $\mathbf{Z}$.
We conducted a small simulation study to compare the performance of different methods to estimate $d$ based on applying the lasso SIR estimator of $y$ on either $\mathcal{Z}$, $\mathcal{X}$, or $\widehat{\mathcal{X}}$. Specifically, after computing the lasso estimator on each dimension $k=1,\ldots,d$, we computed the adjusted eigenvalues as in Lin et al., (2019) and applied $K$-means clustering with $K=2$ on these adjusted eigenvalues. Finally, we estimated $\hat{d}$ as the size of the cluster containing the largest adjusted eigenvalues. To reduce the sensitivity of the final results to the choice of other tuning parameters, we performed $50$ repeated $5$-fold cross-validation (i.e. 50 random splits of the data) where within each repeat, we obtained the lasso SIR estimator for each dimension at the tuning parameter selected as in the simulation, performed the clustering, and recorded the estimated $d$. The final estimator is then recorded as the value of $d$ selected with the highest frequency across 50 repeats. We simulated data using the same setup as in Section 5, but only considered settings with continuous instruments $\mathbf{Z}$ for brevity.
$p$ $n$ $d$ Model Average $\hat{d}$ Proportion $\hat{d}=d$
$\mathbf{Z}$ $\mathbf{X}$ $\widehat{\mathbf{X}}$ $\mathbf{Z}$ $\mathbf{X}$ $\widehat{\mathbf{X}}$
40 200 1 (i) 1.00 1.00 1.00 1.00 1.00 1.00
(ii) 1.00 1.00 1.00 1.00 1.00 1.00
(iii) 1.00 1.00 1.00 1.00 1.00 1.00
2 (iv) 1.76 1.96 1.71 0.75 0.93 0.70
(v) 1.58 1.90 1.52 0.57 0.87 0.49
40 500 1 (i) 1.00 1.00 1.00 1.00 1.00 1.00
(ii) 1.00 1.00 1.00 1.00 1.00 1.00
(iii) 1.00 1.00 1.00 1.00 1.00 1.00
2 (iv) 1.75 1.96 1.72 0.75 0.96 0.72
(v) 1.61 1.92 1.56 0.61 0.92 0.56
500 200 1 (i) 1.00 1.00 1.00 1.00 1.00 1.00
(ii) 1.58 1.11 1.01 0.53 0.89 0.97
(iii) 1.55 1.12 1.02 0.56 0.90 0.97
2 (iv) 2.08 1.77 1.61 0.37 0.54 0.57
(v) 2.10 1.77 1.40 0.41 0.55 0.36
1000 200 1 (i) 1.02 1.00 1.00 0.98 1.00 1.00
(ii) 1.75 1.23 1.09 0.44 0.80 0.91
(iii) 1.68 1.10 1.03 0.52 0.91 0.96
2 (iv) 2.22 1.77 1.68 0.37 0.60 0.36
(v) 2.10 1.87 1.50 0.41 0.45 0.36
Table 3: The mean estimated dimensions from three different lasso-SIR estimators of $y$ on either the instrumental variables $\mathbf{Z}$, the covariates subject to endogeneity $\mathbf{X}$, or the first stage fitted values $\widehat{\mathbf{X}}$. Standard errors are included in parentheses.
Table 3 shows that in settings with $p=40$ and $n\in\{200,500\}$, all three methods performed well in selecting $d=1$ for the three single index models (i)–(iii), while the estimator as applied to the endogenous covariates $\mathbf{X}$ surprisingly had the best performance for the double index models (iv)-(v). Turning to the high-dimensional settings with $n=200$ and $p\in\{500,1000\}$, all three methods still performed well for the linear model (i), but for the non-linear model (ii)-(iii), regressing $y$ on $\mathbf{Z}$ and on $\mathbf{X}$ tended to over-estimate $d$. On the other hand, for the double index models (iv)-(v), the lasso SIR estimator on the instrumental variables $\mathbf{Z}$ performed best while the other two approaches tended to underestimate the dimension.